OpenStudy (anonymous):

Find the second derivative of the function g(t) = -(4/((t+2)^2))

OpenStudy (amistre64):

-4(t+2)^-2

OpenStudy (amistre64):

g' = 8(t+2) g'' = 8

OpenStudy (amistre64):

ugh...soo close

OpenStudy (anonymous):

thats a fraction

OpenStudy (amistre64):

g' = 8(t+2)^-3 g'' = -24(t+2)^-4

OpenStudy (anonymous):

hi amistre! -4/ (t+2)^2

OpenStudy (amistre64):

\[g''=\frac{-24}{(t+2)^4}\]

OpenStudy (anonymous):

oh nvm i see what u did sorry

OpenStudy (amistre64):

product rule tends to be easier to follow that quotient rule

OpenStudy (anonymous):

oh i used the quotient rule

OpenStudy (anonymous):

for g't i got 8t+16/(t^2+4t+4)^2

OpenStudy (amistre64):

we could test it by taking the -4 aside: \[{1\over(t+2)^2}\iff \frac{(t+2)^2 0 - 2(t+2)}{(t+2)^4} \iff \frac{-2}{(t+2)^3}\]

OpenStudy (amistre64):

but it just follows the product rule of it lol

OpenStudy (anonymous):

for g''(t) i got -16t^4-192t^3-480t^2-384t-1008/ (t^4+8t^3+12t^2+32t+16)

OpenStudy (amistre64):

\[g' = {8 \over (t+2)^3} \iff 8(t+2)^{-3} \] right?

OpenStudy (anonymous):

hmm confused

OpenStudy (anonymous):

so what your saying is that our answers are the same? is my answer correct?

OpenStudy (amistre64):

i dont see your answer as being correct; it looks like you got lost in the details

OpenStudy (amistre64):

\[\frac{8}{(t+2)^3}\] doesnt turn into that

OpenStudy (anonymous):

i'll just use the product rule

OpenStudy (amistre64):

\[g'' = {-8.3(t+2)^2 \over (t+2)^6} \iff {-24 \over (t+2)^4}\]

OpenStudy (anonymous):

because it can get pretty contangeled like this

OpenStudy (amistre64):

product rule is simplest :)

OpenStudy (anonymous):

thank you so much :)

OpenStudy (amistre64):

yw :)